OpenStudy (anonymous):

to the nearest whole number. what is the surface area of a sphere with a volume of 140 cm^3

OpenStudy (anonymous):

:-) You should be a wiz at this now.

OpenStudy (anonymous):

lol should be! can be! aint be though lol

OpenStudy (anonymous):

You know that 3V = 4 pi r^3, right?

OpenStudy (anonymous):

haha ummmmm..........yea?????

OpenStudy (anonymous):

now i do

OpenStudy (anonymous):

So what's r?

OpenStudy (anonymous):

11.1

OpenStudy (sriram):

4/3 *pi r^3=v r=[v*(3/4)(1/pi)]^1/3 surface area=4*pi*r^2

OpenStudy (anonymous):

I'm not checking it, I'll just assume you did it right. Now put r in the formula you had before for SA.

OpenStudy (anonymous):

mmm ok

OpenStudy (anonymous):

ie SA = 3*V/r

OpenStudy (anonymous):

umm yah wrong!!!

OpenStudy (anonymous):

fother mucking volume an surface area!!!!! i hate it

OpenStudy (anonymous):

Just check you calculation, did you use a calculator?

OpenStudy (anonymous):

Is r 3.22?

OpenStudy (anonymous):

So 130.43 for SA? (I could have made a mistake as well).

OpenStudy (anonymous):

the closest answer choice is 130 so i think your right so yeah i hope tis is my last volume ?

OpenStudy (anonymous):

I think your problem is not with the surface area / volume, it seems that u are rushing your calculations or making a mistake with the calculator. Just go slowly, check as you go, it will save you time in the end.

OpenStudy (anonymous):

ok so what is the formula for surface area? SA=???

OpenStudy (anonymous):

SA = 3times the volume divided by the radius = 3V/r

OpenStudy (anonymous):

This assumes you know the radius....

OpenStudy (anonymous):

You understand? You have to find the radius with one formula and then the surface area with another.

OpenStudy (anonymous):

ok an yeah i understand

OpenStudy (anonymous):

Just to make it absolutely clear:- Vol = 4*pi*r^3 over 3 (so if you know the volume, you can calculate the radius).

OpenStudy (anonymous):

oh ok thnx

OpenStudy (anonymous):

hey im findin the surface area of a cylynder now does the formula change?

OpenStudy (anonymous):

Including the top and bottom?

OpenStudy (anonymous):

top bottom and the sides

OpenStudy (anonymous):

Well, it's 2 pi r h (for the side) + 2 pi r^2 (for top and bottom together).

OpenStudy (anonymous):

yeag lol i jus found that out i got it from here i think